OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 – 36 = 0 and x2 = 8x – 12 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (anonymous):

@mathslover @amistre64

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@HELP_ME!

OpenStudy (anonymous):

@jadeybells

OpenStudy (triciaal):

part 1 x^2 - 36 do you recognize the difference of 2 squares (x +6)(x-6)=0 x^2 = 8x-12 x^2 - 8x + 12=0 what are the factors of 12 that add up to -8 lets use -6 and -2 (x-6)(x-2)=0